Curriculum Vitae

Jeffrey M. Lessem

Business Address:

Institute for Behavioral Genetics University of Colorado Campus Box 447 Boulder, CO 80309-0447

Phone: 1-303-492-2843
Fax: 1-303-492-0852

e-mail: Jeff.Lessem@Colorado.EDU
WWW: http://ibgwww.colorado.edu/~lessem/

Home Address:

2031 Grandview Ave. Apt. J Boulder, CO 80302-6552

Degrees

B.A., Psychology, University of Texas, 1993

M.A., Psychology, University of Colorado, 1997

Ph.D., Psychology, University of Colorado, 1999

Additional education

1994 Sixth International Workshop on Methodology of Twin and Family Studies, Boulder, CO

1995 Seventh International Workshop on Methodology of Twin and Family Studies, Boulder, CO

1999 Twelfth International Workshop on Methodology of Twin and Family Studies, Boulder, CO (Advanced Course)

Experience

Teaching Assistant

Fall 1995, Introduction to Behavioral Genetics Fall 1997, Concepts in Behavioral Genetics (Computer Skills in Behavior Genetics)

Computer Skills

Systems: Unix (AIX, BSD, Compaq Tru64 (formerly Digital Unix), Linux (Debian, Red

Hat, Slackware), OSF/1, Solaris, System V, Ultrix), OS/2, DOS, Windows 95/98/NT Languages: Fortran, HTML, Perl, shell scripting Networks: TCP/IP, SMB Software: Apache Web Server, Sendmail, Samba, X Window System, Emacs, TEX, L^{A}TEX,

L^{A}TEX 2ε, MySQL, WordPerfect, MS Office Statistical analysis packages: SAS, Mx

Publications

Papers in Refereed Journals

Lessem, J. M., J. K. Hewitt, L. J. Eaves, J. L. Silberg, M. Rutter, E. Simonoff, and J. M. Meyer (1994). Sex Differences in the Occurrence and Rating of Specific Life Events. Behavior Genetics. 24:519-20. (Abstract)

Lessem, J. M., J. K. Hewitt, L. J. Eaves, J. L. Silberg, M. Rutter, E. Simonoff, and J. M. Meyer (1995). Discrepancies in Life Event Reports. Behavior Genetics. 25:274. (Abstract)

Lessem, J. M., J. K. Hewitt, L. J. Eaves, J. L. Silberg, M. Rutter, E. Simonoff, and J. M. Meyer (1996). Life Events and Symptoms of Depression in Adolescents. Behavior Genetics. 26:590. (Abstract)

Lessem, J. M., J. K. Hewitt, L. J. Eaves, J. L. Silberg, M. Rutter, E. Simonoff, and J. M. Meyer (1997). Longitudinal study of life events and depression in adolescence. Behavior Genetics. 27:598. (Abstract)

Lessem, J. M., J. K. Hewitt, L. J. Eaves, J. L. Silberg, M. Rutter, E. Simonoff, and J. M. Meyer (1998). Discrepancies in life event reports with multiple raters and multiple time points. Behavior Genetics. 28:475. (Abstract)

Papers at Refereed Conferences

24^{th} Annual Meeting of the Behavior Genetics Association, Barcelona, Catalonia, Spain, July 1994; Lessem, J. M., J. K. Hewitt, L. J. Eaves, J. L. Silberg, M. Rutter, E. Simonoff, and J. M. Meyer. Sex Differences in the Occurrence and Rating of Specific Life Events. (Poster)

25^{th}Annual Meeting of the Behavior Genetics Association, Richmond, Virginia, USA, May 1995; Lessem, J. M., J. K. Hewitt, L. J. Eaves, J. L. Silberg, M. Rutter, E. Simonoff, and J. M. Meyer. Discrepancies in Life Event Reports. (Talk)

26^{th}Annual Meeting of the Behavior Genetics Association, Pittsburgh, Pennsylvania, USA, June 1996; Lessem, J. M., J. K. Hewitt, L. J. Eaves, J. L. Silberg, M. Rutter, E. Simonoff, and J. M. Meyer. Life Events and Symptoms of Depression in Adolescents. (Poster)

27^{th}Annual Meeting of the Behavior Genetics Association, Toronto, Ontario, Canada, July 1997; Lessem, J. M., J. K. Hewitt, L. J. Eaves, J. L. Silberg, M. Rutter, E. Simonoff, and J. M. Meyer (1997). Longitudinal study of life events and depression in adolescence. (Talk)

27^{th} Annual Meeting of the Behavior Genetics Association, Stockholm, Sweden, June 1998; Lessem, J. M., J. K. Hewitt, L. J. Eaves, J. L. Silberg, M. Rutter, E. Simonoff, and J. M. Meyer (1998). Discrepancies in life event reports with multiple raters and multiple time points. (Talk)

References

Dr. John K. Hewitt University of Colorado Institute for Behavioral Genetics Campus Box 447 Boulder, CO 80309–0447

Dr. John C. DeFries University of Colorado Institute for Behavioral Genetics Campus Box 447 Boulder, CO 80309–0447

November, 1999